# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v4

@@ 3.3.5 Inpainting
Visual inpainting is the process of filling in lost or damaged parts of an image or video. CNNs and GANs have already achieved impressive results (see our previous review paper [8]). Recent work has focused more on editing rather than simply filling in the missing areas. This means users can now mask large areas of an image, and AI tools generate multiple results for users to choose from, a technique known as pluralistic inpainting [300]. Some notable methods include the following: Mask-Aware Transformer (MAT) [149] offers several outputs to fill a large missing area, consisting of a convolutional head, a transformer body, and a convolutional tail for reconstruction, along with a Conv-U-Net for refinement. PUT [150] proposes a patch-based vector VQ-VAE and unquantized Transformer to minimize information loss. Spa-former [153] employs a UNet-like architecture, where each level performs transformer with sparse self-attention to remove coefficients with low or no correlation, leading to memory reduction, while improving result quality by up to 5% compared to PUT.
Video inpainting presents greater complexity compared to image inpainting, despite the abundance of information available in an image sequence. The process typically involves tracking masks across frames, estimating optical flow, and ensuring temporal consistency. The current state-of-the-art methods include DLFormer [151] and ProPainter [152]. DLFormer conducts inpainting in latent space and utilizes discrete codes for video representation. On the other hand, ProPainter employs flow-based deformable alignment to enhance robustness to occlusion and inaccurate flow completion. The method excels in filling complete and rich textures, achieving a speed of 12 fps for full HD video. Video inpainting is also used for dubbing. DINet [301] replaces the mouth area to synchronize with a new language being spoken.
A comprehensive survey of learning-based image and video inpainting, covering approaches such as CNNs, VAEs, GANs, transformers, and diffusion models, can be found in [302]. Additionally, Elharrouss et al. [303] provide an in-depth review of the current challenges and future directions specific to transformer-based inpainting techniques.
@@ 3.3.6 Image Fusion
Transformers and CNNs have been combined to extract global and local information, respectively. Most methods use CNNs for feature extraction, with transformers operating in the latent space [154, 155]. Notable methods include SwinFusion [154], which utilizes a self-attention-based intra-domain fusion unit and a cross-attention-based inter-domain fusion unit to achieve multi-modal and digital photography image fusion. Transformer-based image fusion has also been applied to downstream tasks like segmentation [156], achieving superior results by leveraging the additional information. Self-attention blocks are employed to enhance intra-feature representations, while the cross-attention mechanism integrates inter-feature information to improve the quality of the fused output [157].
DDFM, the first diffusion model-based image fusion method, estimates noise in the reverse process by combining multiple inputs [158]. The expectation-maximization (EM) algorithm is integrated to estimate the noise distribution parameters, resulting in sharper images compared to traditional DDPM. For an in-depth review, the reader is referred to recent work in [304, 305].
@@ 3.3.7 Editing and Visual Special Effects (VFX)
Editing or modifying specific areas of an image is much easier with DM technologies, particularly for headshot photos, such as targeting the eyes and mouth on the face [160]. This capability has been extended to video generation (see Section 3.1.4). Fig. 5 shows an example of the online tool, FacePoke [^44] https://huggingface.co/spaces/jbilcke-hf/FacePoke, which allows users to move the head and modify the shapes of the eyes and mouth in real time. Motion-I2V [159] provides motion blur and motion drag tools to control specific areas of an image to add motion. The method is based on a diffusion-based motion field predictor and motion-augmented temporal attention.
@@
AI technologies can upscale, enhance, and restore low-quality or old footage. For example, standard definition videos can be converted to high definition or even 4K quality without traditional manual remastering processes. This is particularly useful for remastering old movies or enhancing visual details in scenes. Generative AI has also simplified and accelerated automated processes, such as rotoscoping [306], an animation technique where animators trace over motion picture footage frame by frame to create realistic action. AI models can accurately detect and segment objects and characters in video frames, significantly speeding up the post-production process. Additionally, AI can assist the rapid creation of 3D models from 2D images generating realistic animations with minimal input data, facilitating complex human motions and synchronized facial expressions to voiceovers. One restriction is that current technologies still cannot yet generate full 4K accurate visual effects.
@@ 3.4 Information Extraction and Understanding
AI plays a crucial role in automating and optimizing the process of information extraction and understanding, enabling organizations to derive actionable insights from large and diverse data. Yan et al. [91] have categorized information extraction tasks based on the Format-Time-Reference space, as illustrated in Fig. 6 (a), where object detection and video object segmentation (VOS) are considered to be the simplest and the most complex tasks, respectively. Recent advancements in this field draw significant inspiration from LLMs. These advancements include the utilization of prompts as conditional inputs for acquiring information. Moreover, following the pipeline approach used in LLMs, there is a growing trend towards leveraging very large datasets to pre-train models before fine-tuning them for specific downstream tasks. For instance, Meta AI [166] has introduced DINOv2, aimed at enriching information about visual content through self-supervised learning. This model was trained with 142 million carefully selected images, employing the ViT architecture. Google have introduced VideoPrism [178], a tool for scene understanding including classification, localization, retrieval, captioning, and question answering (QA). The model was trained on an extensive and diverse dataset consisting of 36 million high-quality video-text pairs and 582 million video clips accompanied by noisy or machine-generated parallel text.
Figure 6: (a) Tasks in Object-centric understanding defined by Yan et al. [91] (REC:, Referring Expression Comprehension, RES: Referring Expression Segmentation, VOS: Video Object Segmentation, RVOS: Referring Video Object Segmentation, MOT: Multiple Object Tracking, MOTS: Multi-Object Tracking and Segmentation, VIS: Video Instance Segmentation, SOT: Single Object Tracking. (b) Current high-quality segmentation [163].
@@ 3.4.1 Segmentation
The need for segmentation has grown dramatically in the past few years, given its central role in visual perception. Many segmentation methods now integrate an input prompt for users to define their preferred output appearances, such as pixel-wise segmentation, bounding boxes around objects, or segmented areas of interest. Most of these methods utilize transformer architectures [161]. Among them, Segment Anything (SAM) by Meta AI [162] stands out as a pioneer in promptable segmentation approaches. This method computes masks in real-time and has been trained with over 1 billion masks across 11 million images, facilitating transferability from zero-shot to new image distributions and tasks. HQ-SAM [163] enhances SAM by incorporating global-local feature fusion, leading to high-quality mask predictions. SegGPT [164] proposed context ensemble strategies and allows users to tune a prompt for a specific dataset, scene, or even a person, while SEEM [165] provides a completely promptable and interactive segmentation interface. More recently, SAM2 [167] introduced support for real-time video segmentation. It is a unified model trained on a larger dataset than SAM. Interactive tools enable users to mark areas of interest and specify regions to exclude from the segmentation map. Zhou et al. propose an audio-visual segmentation (AVS) to generate pixel-level segmentation masks for sounding objects in audible videos. DVIS++ by [168] introduces a universal video segmentation framework capable of producing instance, semantic, and panoptic segmentation outputs. This transformer-based architecture comprises a segmentor, tracker, and refinement module, achieving state-of-the-art performance across several video segmentation benchmarks.
@@
With DMs tehchnologies, Baranchuk et al. [307] have investigated semantic representation, and found DMs outperform other few-shot learning approaches. DiffuMask [169] automatically generate image and pixel-level semantic annotation using pre-trained Stable Diffusion with input as a text prompt. It has been proven that using these synthetic data improve segmentation accuracy. Currently, the state-of-the-art panoptic segmentation is the method developed by Nvdia, which is based on text-to-image DMs [170], outperforming the previous methods by up to 7.6%.
@@
Applying INRs to segmentation is more popular in the medical domain, as the specific signals used, such as computed tomography (CT) and magnetic resonance imaging (MRI), can be formulated as continuous functions. In creative technologies, unsupervised domain adaptation (UDA) and INRs are used for continuous rectification function modeling in [172], achieving superior segmentation results in night vision. Recently, this work has been integrated with a non-local means block in [308] showning a significant improvement for instant segmentation in low-light scenes.
@@
3D segmentation is also crucial for scene manipulation. In radiance fields, earlier segmentation methods required additional modules such as using k-means clustering to separate objects from the background [309]. However, the recent SA3D approach [173] segments 3D objects using NeRFs as the structural prior. SA3D operates by taking a trained NeRF and a set of prompts from a single view, then performing an iterative procedure. This involves rendering novel 2D views, self-prompting SAM for 2D segmentation, and projecting the segmentation back onto 3D mask grids. A comprehensive survey of 3D segmentation in computer vision can be found in [310].
@@ 3.4.2 Detection and recognition
Introduced in 2020, DETR by Facebook AI [174] was one of the first to adopt a transformer architecture for object detection. The approach achieves comparable results to an optimized Faster R-CNN [311], introduced in 2015. Deformable convolution has alson been used, (Deformable DETR [175]), resulting in training faster with approximately 5% accuracy improvement. A survey until 2022 [312] reported that Deformable DETR and Swin Transformers [16] outperform pure CNN-based YOLOv4 [313]. SwinV2 improves the first version by replacing original dot product attention with scaled cosine attention, improving accuracy by approximately 5%. Later, RT-DETR [314] improved inference speed by decoupling the intra-scale interaction and cross-scale fusion of features with different scales. RT-DETR is 25% faster than YOLOv8464646https://github.com/ultralytics/ultralytics with 6% improvement on MS COCO Object Detection dataset. Recently, YOLOv10 [315] has been released. YOLOv10 further improves the speed of detection approximately by 30% over RT-DETR with the same accuracy. A review of transformer-based methods for object detection can be found in [316, 317]. Recently, YOLO12 [180] introduced an attention-centric architecture, achieving a 2.1% and 1.2% mAP improvement over YOLOv10-N and YOLOv11-N respectively, with only a slight decrease in speed.
To detect 3D objects, the transformer-based method MonoDTR [177] incorporates depth estimation from a single 2D image [201] to predict 3D bounding boxes. More 3D object detection methods have been developed for autonomous driving [318]; however, these approaches can also be adapted for AR and VR applications [179].
While DMs are primarily used to generate synthetic datasets [319, 320], they have also been demonstrated to function as zero-shot classifiers by Li et al. [181]. DMs are also of interest for detection tasks, Although feature extractors are still predominantly based on CNNs, such as ResNet, or Transformers (like Swin). DiffusionDet [182] formulates object detection as a denoising diffusion process from noisy boxes to object boxes, reporting performance that surpasses DETR. DMs have also been employed for anomaly detection [183, 184], functioning similarly to zero-shot classifiers.
@@ 3.4.3 Tracking
Object tracking stands out as one of the tasks that greatly benefits from transformers since attention is needed in both space and time. An experimental survey cited in [321] reveals that transformer-based methods consistently rank at the top of the leaderboard across various datasets. In the Visual Object Tracking (VOT) challenges of 2023474747https://eu.aihub.ml/competitions/201#results, all of the top-10 employed transformer-based methodologies. The highest-performing approach achieved a 10% improvement in tracking quality compared to the winner in 2020. The current state-of-the-art for single-object tracking [^48] https://paperswithcode.com/sota/visual-object-tracking-on-lasot, however, is based on cross-attention and Mamba [193].
The first three tracking-by-attention approaches are TrackFormer [185], MixFormer [187], and ToMP [188]. TrackFormer extracts visual features using a CNN-based encoder, which are then tracked using a vanilla transformer [10] in a frame sequence, while MixFormer introduces cross-attention between the target and search regions. ToMP tracks the objects using prediction aspects. Many more methods have been proposed, including SeqTrack [190] and Track Anything Model (TAM) [189]. SeqTrack extracts visual features with a bidirectional transformer, while the decoder generates a sequence of bounding box values autoregressively with a causal transformer. TAM combines SAM [162] and XMem [322], offering tracking and segmentation performance on the human-selected target. However, the masked area is still not very sharp, and there is a subtle degree of temporal inconsistency. MOTRv2 [191] combines YOLOX [323] for object recognition and MOTR [186] for tracking, outperforming TrackFormer by 20%. Additionally, some methods have been specifically proposed for challenging environments, such as low light [192] and small objects, as seen in AnyFlow [197]. The latter exploits INR to upsample a continuous coordinate-based flow map, similar to SISR technique proposed in [127].
Similarly to detection tasks, DMs for tracking tasks are used as downstream processes by concatenating the diffusion head to the feature extraction backbone. However, a spatial-temporal fusion module has been added to the diffusion head to exploit temporal video features[194]. DiffusionTrack [195] localizes the target in a progressive diffusion manner, which is claimed to better handle challenging scenarios. The method in [196] exploits spatial-temporal weighting to suppress the probability of the tracker changing the target to the distractors. It, however, reports underperformance compared to MixFormer.

### v5

@@ 3.3.5 Inpainting
Visual inpainting is the process of filling in lost or damaged parts of an image or video. CNNs and GANs have already achieved impressive results (see our previous review paper [Anantrasirichai:AI:2022]). Recent work has focused more on editing rather than simply filling in the missing areas. This means users can now mask large areas of an image, and AI tools generate multiple results for users to choose from, a technique known as pluralistic inpainting [zheng:pluralistic:2019]. Some notable methods include the following: Mask-Aware Transformer (MAT) [Li:MAT:2022] offers several outputs to fill a large missing area, consisting of a convolutional head, a transformer body, and a convolutional tail for reconstruction, along with a Conv-U-Net for refinement. PUT [Liu:Reduce:2022] proposes a patch-based vector VQ-VAE and unquantized Transformer to minimize information loss. Spa-former [HUANG:Sparse:2024] employs a UNet-like architecture, where each level performs transformer with sparse self-attention to remove coefficients with low or no correlation, leading to memory reduction, while improving result quality by up to 5% compared to PUT.
Video inpainting presents greater complexity compared to image inpainting, despite the abundance of information available in an image sequence. The process typically involves tracking masks across frames, estimating optical flow, and ensuring temporal consistency. The current state-of-the-art methods include DLFormer [Ren:DLFormer:2022] and ProPainter [Zhou:ProPainter:2023]. DLFormer conducts inpainting in latent space and utilizes discrete codes for video representation. On the other hand, ProPainter employs flow-based deformable alignment to enhance robustness to occlusion and inaccurate flow completion. The method excels in filling complete and rich textures, achieving a speed of 12 fps for full HD video. Video inpainting is also used for dubbing. DINet [Zhang:DINet:2023] replaces the mouth area to synchronize with a new language being spoken.
A comprehensive survey of learning-based image and video inpainting, covering approaches such as CNNs, VAEs, GANs, transformers, and diffusion models, can be found in [quan:deep:2024]. Additionally, Elharrouss et al. [elharrouss2025transformer] provide an in-depth review of the current challenges and future directions specific to transformer-based inpainting techniques.
@@ 3.3.6 Image Fusion
Transformers and CNNs have been combined to extract global and local information, respectively. Most methods use CNNs for feature extraction, with transformers operating in the latent space [Ma:SwinFusion:2022, Rao:TGFuse:2023]. Notable methods include SwinFusion [Ma:SwinFusion:2022], which utilizes a self-attention-based intra-domain fusion unit and a cross-attention-based inter-domain fusion unit to achieve multi-modal and digital photography image fusion. Transformer-based image fusion has also been applied to downstream tasks like segmentation [Liu:Multi:2023], achieving superior results by leveraging the additional information. Self-attention blocks are employed to enhance intra-feature representations, while the cross-attention mechanism integrates inter-feature information to improve the quality of the fused output [LI2024102147].
DDFM, the first diffusion model-based image fusion method, estimates noise in the reverse process by combining multiple inputs [Zhao:DDFM:2023]. The expectation-maximization (EM) algorithm is integrated to estimate the noise distribution parameters, resulting in sharper images compared to traditional DDPM. For an in-depth review, the reader is referred to recent work in [Karim:Current:2023, Zhang:Visible:2023].
@@ 3.3.7 Editing and Visual Special Effects (VFX)
Editing or modifying specific areas of an image is much easier with DM technologies, particularly for headshot photos, such as targeting the eyes and mouth on the face [guo2024liveportrait]. This capability has been extended to video generation (see Section 3.1.4). Fig. 5 shows an example of the online tool, FacePoke [^44] https://huggingface.co/spaces/jbilcke-hf/FacePoke, which allows users to move the head and modify the shapes of the eyes and mouth in real time. Motion-I2V [Shi:Motion-I2V:2024] provides motion blur and motion drag tools to control specific areas of an image to add motion. The method is based on a diffusion-based motion field predictor and motion-augmented temporal attention.
@@
AI technologies can upscale, enhance, and restore low-quality or old footage. For example, standard definition videos can be converted to high definition or even 4K quality without traditional manual remastering processes. This is particularly useful for remastering old movies or enhancing visual details in scenes. Generative AI has also simplified and accelerated automated processes, such as rotoscoping [Tous:Lester:2024], an animation technique where animators trace over motion picture footage frame by frame to create realistic action. AI models can accurately detect and segment objects and characters in video frames, significantly speeding up the post-production process. Additionally, AI can assist the rapid creation of 3D models from 2D images generating realistic animations with minimal input data, facilitating complex human motions and synchronized facial expressions to voiceovers. One restriction is that current technologies still cannot yet generate full 4K accurate visual effects.
@@ 3.4 Information Extraction and Understanding
AI plays a crucial role in automating and optimizing the process of information extraction and understanding, enabling organizations to derive actionable insights from large and diverse data. Yan et al. [Yan:Universal:2023] have categorized information extraction tasks based on the Format-Time-Reference space, as illustrated in Fig. 6 (a), where object detection and video object segmentation (VOS) are considered to be the simplest and the most complex tasks, respectively. Recent advancements in this field draw significant inspiration from LLMs. These advancements include the utilization of prompts as conditional inputs for acquiring information. Moreover, following the pipeline approach used in LLMs, there is a growing trend towards leveraging very large datasets to pre-train models before fine-tuning them for specific downstream tasks. For instance, Meta AI [Oquab:DINOv2:2024] has introduced DINOv2, aimed at enriching information about visual content through self-supervised learning. This model was trained with 142 million carefully selected images, employing the ViT architecture. Google have introduced VideoPrism [zhao:videoprism:2024], a tool for scene understanding including classification, localization, retrieval, captioning, and question answering (QA). The model was trained on an extensive and diverse dataset consisting of 36 million high-quality video-text pairs and 582 million video clips accompanied by noisy or machine-generated parallel text.
Figure 6: (a) Tasks in Object-centric understanding defined by Yan et al. [Yan:Universal:2023] (REC:, Referring Expression Comprehension, RES: Referring Expression Segmentation, VOS: Video Object Segmentation, RVOS: Referring Video Object Segmentation, MOT: Multiple Object Tracking, MOTS: Multi-Object Tracking and Segmentation, VIS: Video Instance Segmentation, SOT: Single Object Tracking. (b) Current high-quality segmentation [Ke:SAM-HQ:2023].
@@ 3.4.1 Segmentation
The need for segmentation has grown dramatically in the past few years, given its central role in visual perception. Many segmentation methods now integrate an input prompt for users to define their preferred output appearances, such as pixel-wise segmentation, bounding boxes around objects, or segmented areas of interest. Most of these methods utilize transformer architectures [Bowen:Marked:2022]. Among them, Segment Anything (SAM) by Meta AI [Kirillov:SAM:2023] stands out as a pioneer in promptable segmentation approaches. This method computes masks in real-time and has been trained with over 1 billion masks across 11 million images, facilitating transferability from zero-shot to new image distributions and tasks. HQ-SAM [Ke:SAM-HQ:2023] enhances SAM by incorporating global-local feature fusion, leading to high-quality mask predictions. SegGPT [Wang:SegGPT:2023] proposed context ensemble strategies and allows users to tune a prompt for a specific dataset, scene, or even a person, while SEEM [Zou:Segment:2023] provides a completely promptable and interactive segmentation interface. More recently, SAM2 [ravi2024sam2] introduced support for real-time video segmentation. It is a unified model trained on a larger dataset than SAM. Interactive tools enable users to mark areas of interest and specify regions to exclude from the segmentation map. Zhou et al. propose an audio-visual segmentation (AVS) to generate pixel-level segmentation masks for sounding objects in audible videos. DVIS++ by [Zhang:DVISp:2025] introduces a universal video segmentation framework capable of producing instance, semantic, and panoptic segmentation outputs. This transformer-based architecture comprises a segmentor, tracker, and refinement module, achieving state-of-the-art performance across several video segmentation benchmarks.
@@
With DMs tehchnologies, Baranchuk et al. [Baranchuk:label:2022] have investigated semantic representation, and found DMs outperform other few-shot learning approaches. DiffuMask [Wu:DiffuMask:2023] automatically generate image and pixel-level semantic annotation using pre-trained Stable Diffusion with input as a text prompt. It has been proven that using these synthetic data improve segmentation accuracy. Currently, the state-of-the-art panoptic segmentation is the method developed by Nvdia, which is based on text-to-image DMs [Xu:Open:2023], outperforming the previous methods by up to 7.6%.
@@
Applying INRs to segmentation is more popular in the medical domain, as the specific signals used, such as computed tomography (CT) and magnetic resonance imaging (MRI), can be formulated as continuous functions. In creative technologies, unsupervised domain adaptation (UDA) and INRs are used for continuous rectification function modeling in [Gong:Continuous:2023], achieving superior segmentation results in night vision. Recently, this work has been integrated with a non-local means block in [Lin:Feature:2024] showning a significant improvement for instant segmentation in low-light scenes.
@@
3D segmentation is also crucial for scene manipulation. In radiance fields, earlier segmentation methods required additional modules such as using k-means clustering to separate objects from the background [Goel:Interactive:2023]. However, the recent SA3D approach [Cen:Segment:2023] segments 3D objects using NeRFs as the structural prior. SA3D operates by taking a trained NeRF and a set of prompts from a single view, then performing an iterative procedure. This involves rendering novel 2D views, self-prompting SAM for 2D segmentation, and projecting the segmentation back onto 3D mask grids. A comprehensive survey of 3D segmentation in computer vision can be found in [HE2025102722].
@@ 3.4.2 Detection and recognition
Introduced in 2020, DETR by Facebook AI [Carion:DERT:2020] was one of the first to adopt a transformer architecture for object detection. The approach achieves comparable results to an optimized Faster R-CNN [Ren:Faster:2027], introduced in 2015. Deformable convolution has alson been used, (Deformable DETR [Zhu:Deformable:2021]), resulting in training faster with approximately 5% accuracy improvement. A survey until 2022 [Zou:object:2023] reported that Deformable DETR and Swin Transformers [Liu:Swin:2021] outperform pure CNN-based YOLOv4 [bochkovskiy2020yolov4]. SwinV2 improves the first version by replacing original dot product attention with scaled cosine attention, improving accuracy by approximately 5%. Later, RT-DETR [lv2:detrs:2024] improved inference speed by decoupling the intra-scale interaction and cross-scale fusion of features with different scales. RT-DETR is 25% faster than YOLOv8464646https://github.com/ultralytics/ultralytics with 6% improvement on MS COCO Object Detection dataset. Recently, YOLOv10 [wang:yolov10:2024] has been released. YOLOv10 further improves the speed of detection approximately by 30% over RT-DETR with the same accuracy. A review of transformer-based methods for object detection can be found in [Li:Transformer:2023, KHEDDAR2025103347]. Recently, YOLO12 [tian2025yolov12] introduced an attention-centric architecture, achieving a 2.1% and 1.2% mAP improvement over YOLOv10-N and YOLOv11-N respectively, with only a slight decrease in speed.
To detect 3D objects, the transformer-based method MonoDTR [Huang:MonoDTR:2022] incorporates depth estimation from a single 2D image [Yang:depthanything:2024] to predict 3D bounding boxes. More 3D object detection methods have been developed for autonomous driving [10637966]; however, these approaches can also be adapted for AR and VR applications [im2025gate3d].
While DMs are primarily used to generate synthetic datasets [Wu:datasetDM:2023, Fang:Data:2024], they have also been demonstrated to function as zero-shot classifiers by Li et al. [Li:Your:2023]. DMs are also of interest for detection tasks, Although feature extractors are still predominantly based on CNNs, such as ResNet, or Transformers (like Swin). DiffusionDet [Chen:DiffusionDet:2023] formulates object detection as a denoising diffusion process from noisy boxes to object boxes, reporting performance that surpasses DETR. DMs have also been employed for anomaly detection [Zhang:DiffAD:2025, WU2025102965], functioning similarly to zero-shot classifiers.
@@ 3.4.3 Tracking
Object tracking stands out as one of the tasks that greatly benefits from transformers since attention is needed in both space and time. An experimental survey cited in [Kugarajeevan:Transformers:2023] reveals that transformer-based methods consistently rank at the top of the leaderboard across various datasets. In the Visual Object Tracking (VOT) challenges of 2023474747https://eu.aihub.ml/competitions/201#results, all of the top-10 employed transformer-based methodologies. The highest-performing approach achieved a 10% improvement in tracking quality compared to the winner in 2020. The current state-of-the-art for single-object tracking [^48] https://paperswithcode.com/sota/visual-object-tracking-on-lasot, however, is based on cross-attention and Mamba [kang2025exploring].
The first three tracking-by-attention approaches are TrackFormer [Meinhardt:TrackFormer:2022], MixFormer [cui:mixformer:2022], and ToMP [Mayer:Transforming:2022]. TrackFormer extracts visual features using a CNN-based encoder, which are then tracked using a vanilla transformer [Vaswani:attention:2017] in a frame sequence, while MixFormer introduces cross-attention between the target and search regions. ToMP tracks the objects using prediction aspects. Many more methods have been proposed, including SeqTrack [Chen:SeqTrack:2023] and Track Anything Model (TAM) [yang:track:2023]. SeqTrack extracts visual features with a bidirectional transformer, while the decoder generates a sequence of bounding box values autoregressively with a causal transformer. TAM combines SAM [Kirillov:SAM:2023] and XMem [cheng:xmem:2022], offering tracking and segmentation performance on the human-selected target. However, the masked area is still not very sharp, and there is a subtle degree of temporal inconsistency. MOTRv2 [Zhang:MOTRv2:2023] combines YOLOX [ge:yolox:2021] for object recognition and MOTR [zeng:motr:2022] for tracking, outperforming TrackFormer by 20%. Additionally, some methods have been specifically proposed for challenging environments, such as low light [Yi:Comprehensive:2024] and small objects, as seen in AnyFlow [Jung:AnyFlow:2023]. The latter exploits INR to upsample a continuous coordinate-based flow map, similar to SISR technique proposed in [Chen:Learning:2021].
Similarly to detection tasks, DMs for tracking tasks are used as downstream processes by concatenating the diffusion head to the feature extraction backbone. However, a spatial-temporal fusion module has been added to the diffusion head to exploit temporal video features[Luo:DiffusionTrack:2024]. DiffusionTrack [Xie:DiffusionTrack:2024] localizes the target in a progressive diffusion manner, which is claimed to better handle challenging scenarios. The method in [Zhang:DiffusionTracker:2024] exploits spatial-temporal weighting to suppress the probability of the tracker changing the target to the distractors. It, however, reports underperformance compared to MixFormer.